Which of the following statements are wrong?
Orbital angular momentum of 1s,2s and 3s electrons are same
Mn+7 is diamagnetic
Spin magnetic moment of Fe(0) is 6.92 B.M.
Percentage uncertainty in velocity of an electron in motion is equal to percentage uncertainty in its wavelength
1
Expert's answer
2020-06-12T07:29:55-0400

The permanganate ion (Mn+7) is diamagnetic since it has a tetrahedral structure with no unpaired electrons.

In 1s,2s and 3s orbitals , the orbital angular momentum of the electrons is the same since they have an equal of azimuthal quantum number .

"\mu n=\sqrt{ l(l+n)}\dfrac{h}{2\pi}" where for all s orbitals l=0

FeO has 6 unpaired electrons.

Spin magnetic moment="\sqrt{6(6+2)}=6.97B.M"

The last statement is not true because percentage of uncertainty in velocity of an electron in motion is not equal to the percentage uncertainty in its wavelength.